# ASDF: Assembly State Detection Utilizing Late Fusion by Integrating 6D Pose Estimation

### v2

@@ 4 Discussion
ASDF uses deep learning-based 6D position estimation and assembly state detection. Our approach shows more robust results compared to a pure deep learning-based approach on our ASDF and the GBOT dataset. Robust 6D pose estimation and assembly state detection allow adding visual overlays, e.g. see Figure 7 or even highlighting the current assembly state , see Figure 4. The more robust and correct the prediction is, the more reliably an AR interface can display this information.
@@
The state transition, i.e. the change from state A to B, is a frequent challenge not only in the detection of assembly states. However, it is generally a challenge in phase detection or action detection [6]. A common approach in this regard, which is not considered in our evaluation, is to define a time frame between transitions, which is not considered in the evaluation. We did not do this as we were aiming for a real-world scenario where constant knowledge of the assembly state and object position is crucial. Therefore, we included these potentially error-prone parts in our dataset. As shown in Figure 8, ASDF is challenged by this transition, but can retrieve more robust results with our Pose2State module compared to the pure deep learning-based approach.
@@ 4.1 Limitations
The current approach focuses on reliable pose estimations and assembly state detection using a single camera. To address occlusion a multi-camera approach or even fusing static and dynamic camera input could provide additional information for the estimation.
@@ 5 Conclusion
ASDF is a deep learning-based approach that utilizes late fusion of assembly state and 6D pose estimation to enhance its prediction. Our approach enables smart guidance as it can continuously track the 6D object pose and state of several assembly objects. This can be utilized in the medical or industry context. ASDF can track objects from assembly state zero (unassembled) until the full assembly is completed. Moreover, it can better detect state changes compared to a pure deep learning-based approach using the Pose2State module.
In addition to our assembly dataset, where we outperform the pure deep learning-based approach, ASDF has demonstrated that awareness of the assembly state leads to an improved performance compared to the state-of-the-art on the GBOT dataset. The scenes in these datasets present various challenges, and ASDF has shown significant improvement compared to YOLOv8Pose, even outperforming GBOT.
@@ Acknowledgements.
The authors thank Philipp Stefan, Patrick Wucherer, and Matthew Lau from Medability GmbH for providing the 3D printed assembly parts. This work is funded by The German Federal Ministry of Education and Research (BMBF) with grant number 16SV8973. We thank d.hip for providing a campus stipend.References
@@
- [3] S. Borkman, A. Crespi, S. Dhakad, S. Ganguly, J. Hogins, Y.-C. Jhang, M. Kamalzadeh, B. Li, S. Leal, P. Parisi, C. Romero, W. Smith, A. Thaman, S. Warren, and N. Yadav. Unity perception: Generate synthetic data for computer vision, 2021.
- [4] E. Bottani and G. Vignali. Augmented reality technology in the manufacturing industry: A review of the last decade. IISE Transactions, 51(3):284–310, 2019. doi: 10 . 1080/24725854 . 2018 . 1493244
@@
- [8] M. Eswaran, A. K. Gulivindala, A. K. Inkulu, and M. V. A. R. Bahubalendruni. Augmented reality-based guidance in product assembly and maintenance/repair perspective: A state of the art review on challenges and opportunities. Expert Systems with Applications, 2023.
@@
- [13] T. Hodan, J. Matas, and S. Obdrzalek. On evaluation of 6d object pose estimation. vol. 9915, pp. 606–619. Springer International Publishing, 2016. doi: 10 . 1007/978-3-319-49409-8_52
@@
- [15] T. Jantos, M. Hamdad, W. Granig, S. Weiss, and J. Steinbrener. PoET: Pose estimation transformer for single-view, multi-object 6d pose estimation. In 6th Annual Conference on Robot Learning (CoRL 2022), 2022.
@@
- [19] S. Li, H. Schieber, N. Corell, B. Egger, J. Kreimeier, and D. Roth. GBOT: Graph-Based 3D Object Tracking for Augmented Reality-Assisted Assembly Guidance, 2024.
- [20] Y. Li, G. Wang, X. Ji, Y. Xiang, and D. Fox. Deepim: Deep iterative matching for 6d pose estimation. In Proceedings of the European Conference on Computer Vision (ECCV), pp. 683–698, 2018.
- [21] Y. Li, F. Zhong, X. Wang, S. Song, J. Li, X. Qin, and C. Tu. For a more comprehensive evaluation of 6dof object pose tracking. 2023.
@@
- [24] K. Murray, J. Schierl, K. Foley, and Z. Duric. Equipment assembly recognition for augmented reality guidance. pp. 109–118, 2024.
@@
- [26] S. Peng, Y. Liu, Q. Huang, X. Zhou, and H. Bao. Pvnet: Pixel-wise voting network for 6dof pose estimation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 4561–4570, 2019.
- [27] W. Qiu and A. Yuille. UnrealCV: Connecting computer vision to unreal engine. In European Conference on Computer Vision Workshops (ECCVW), pp. 909–916, 2016.
- [28] J. Rambach, A. Pagani, M. Schneider, O. Artemenko, and D. Stricker. 6dof object tracking based on 3d scans for augmented reality remote live support. Computers, 7(1):6, 2018.
@@
- [32] M. Stoiber, M. Elsayed, A. E. Reichert, F. Steidle, D. Lee, and R. Triebel. Fusing visual appearance and geometry for multi-modality 6dof object tracking, 2023. Publication Title: arXiv preprint arXiv:2302.11458.
@@
- [38] B. Tekin, S. N. Sinha, and P. Fua. Real-time seamless single shot 6d object pose prediction. In Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition (CVPR), 2018.
@@
- [41] G. Wang, F. Manhardt, F. Tombari, and X. Ji. Gdr-net: Geometry-guided direct regression network for monocular 6d object pose estimation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 16611–16621, 2021.
- [42] H. Wang, S. Sridhar, J. Huang, J. Valentin, S. Song, and L. J. Guibas. Normalized object coordinate space for category-level 6d object pose and size estimation. In The IEEE Conference on Computer Vision and Pattern Recognition (CVPR), 2019.

### v3

@@ 4 Discussion
ASDF shows more robust results compared to a pure deep learning-based approach on our ASDF and the GBOT dataset. Robust 6D pose estimation and assembly state detection allow adding visual overlays, e.g. see Figure 6 or even highlighting the current assembly state, see Figure 4. The more robust and correct the prediction is, the more reliably an AR interface can display this information.
@@
The state transition, i.e. the change from state A to B, is a frequent challenge not only in the detection of assembly states. However, it is generally a challenge in phase detection or action detection [5]. A common approach in this regard, which is not considered in our evaluation, is to define a time frame between transitions, which is not considered in the evaluation. We did not do this as we were aiming for a real-world scenario where constant knowledge of the assembly state and object position is crucial. Therefore, we included these potentially error-prone parts in our dataset.
@@ 4.1 Limitations
Moreover, the ASDF dataset features texture-less objects which do not propose the challenge of reflective materials such as medical instruments. Nevertheless, texture less objects propose a challenge.
The current approach focuses on reliable pose estimations and assembly state detection using a single camera and for the dataset on reproducible items. To address occlusion a multi-camera approach or even fusing static and dynamic camera input could provide additional information for the estimation.
@@ 5 Conclusion
We present ASDF a deep learning-based approach utilizing late fusion for assembly state detection and 6D pose estimation. ASDF can track objects from assembly state zero (unassembled) until the full assembly is completed. This enables smart guidance and can be used in the medical or industrial context. On our assembly dataset, we outperform the deep learning-based without a fusion step. Moreover, ASDF has demonstrated that awareness of the assembly state leads to an improved performance compared to the state-of-the-art in 6D pose estimation on the GBOT dataset. The scenes in this dataset present various challenges, and ASDF has shown a huge improvement compared to YOLOv8Pose and outperformed all object tracking-based approaches.
@@ Acknowledgements.
This work is funded by the German Federal Ministry of Education and Research (BMBF) with grant number 16SV8973. We further thank d.hip for providing Hannah Schieber with a campus stipend.References
@@
- [4] E. Cramer, A. Kucharski, J. Kreimeier, S. Andreß, S. Li, C. Walk, F. Merkl, J. Högl, P. Wucherer, P. Stefan, et al. Requirement analysis for an ai-based ar assistance system for surgical tools in the operating room: stakeholder requirements and technical perspectives. International Journal of Computer Assisted Radiology and Surgery, pp. 1–10, 2024.
@@
- [7] M. A. Fischler and R. C. Bolles. Random sample consensus: a paradigm for model fitting with applications to image analysis and automated cartography. Communications of the ACM, 24(6):381–395, 1981.
- [8] R. Girshick. Fast r-cnn. In Proceedings of the IEEE international conference on computer vision, pp. 1440–1448. IEEE/CVF, 2015.
@@
- [11] S. J. Henderson and S. K. Feiner. Augmented reality in the psychomotor phase of a procedural task. In 2011 10th IEEE International Symposium on Mixed and Augmented Reality, pp. 191–200, 2011. doi: 10 . 1109/ISMAR . 2011 . 6092386
@@
- [14] T. Hodan, J. Matas, and S. Obdrzalek. On evaluation of 6d object pose estimation. In G. Hua and H. Jegou, eds., Computer Vision–ECCV 2016 Workshops: Amsterdam, The Netherlands, October 8-10 and 15-16, 2016, Proceedings, Part III 14, vol. 9915, pp. 606–619. Springer International Publishing, 2016. doi: 10 . 1007/978-3-319-49409-8_52
@@
- [16] G. Jocher, A. Chaurasia, and J. Qiu. YOLO by ultralytics, 2023.
@@
- [20] V. Lepetit, F. Moreno-Noguer, and P. Fua. Ep n p: An accurate o (n) solution to the p n p problem. International journal of computer vision, 81:155–166, 2009.
- [21] S. Li, H. Schieber, N. Corell, B. Egger, J. Kreimeier, and D. Roth. GBOT: Graph-based 3d object tracking for augmented reality-assisted assembly guidance. In 2024 IEEE Conference Virtual Reality and 3D User Interfaces (VR), pp. 513–523. IEEE, 2024.
- [22] Y. Li, F. Zhong, X. Wang, S. Song, J. Li, X. Qin, and C. Tu. For a more comprehensive evaluation of 6dof object pose tracking, 2023.
@@
- [25] K. Murray, J. Schierl, K. Foley, and Z. Duric. Equipment assembly recognition for augmented reality guidance. In 2024 IEEE International Conference on Artificial Intelligence and eXtended and Virtual Reality (AIxVR), pp. 109–118. IEEE, 2024.
@@
- [27] S. Peng, Y. Liu, Q. Huang, X. Zhou, and H. Bao. Pvnet: Pixel-wise voting network for 6dof pose estimation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 4561–4570. IEEE/CVF, 2019.
- [28] R. Radkowski. Object tracking with a range camera for augmented reality assembly assistance. Journal of Computing and Information Science in Engineering, 16(1):011004, 2016.
@@
- [31] T. J. Schoonbeek, T. Houben, H. Onvlee, P. H. N. d. With, and F. v. d. Sommen. IndustReal: A dataset for procedure step recognition handling execution errors in egocentric videos in an industrial-like setting, 2024.
@@
- [33] M. Stoiber, M. Elsayed, A. E. Reichert, F. Steidle, D. Lee, and R. Triebel. Fusing visual appearance and geometry for multi-modality 6dof object tracking. In 2023 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS), pp. 1170–1177. IEEE, 2023.
@@
- [39] A. Tang, C. Owen, F. Biocca, and W. Mou. Comparative effectiveness of augmented reality in object assembly. In Proceedings of the SIGCHI conference on Human factors in computing systems, pp. 73–80, 2003.
- [40] B. Tekin, S. N. Sinha, and P. Fua. Real-time seamless single shot 6d object pose prediction. In Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition (CVPR). IEEE/CVF, 2018.
- [41] K. K. Thiel. Automated Creation, Evaluation and Configuration of Markerless Object Tracking for Superimposed Augmented Reality. PhD thesis, Technische Universität München, 2023.
@@
- [44] VisionLib. Showcase: Tracking multiple parts - visionlib. https://www.youtube.com/watch?v=GfO6hmk7kww, 2024. Accessed: 2024-06-26.
- [45] G. Wang, F. Manhardt, F. Tombari, and X. Ji. Gdr-net: Geometry-guided direct regression network for monocular 6d object pose estimation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 16611–16621. IEEE/CVF, 2021.
@@
- [50] M. Zaccaria, F. Manhardt, Y. Di, F. Tombari, J. Aleotti, and M. Giorgini. Self-supervised category-level 6d object pose estimation with optical flow consistency. IEEE Robotics and Automation Letters, 8(5):2510–2517, 2023. doi: 10 . 1109/LRA . 2023 . 3254463
@@
- [52] Y. Zheng, Y. Kuang, S. Sugimoto, K. Astrom, and M. Okutomi. Revisiting the pnp problem: A fast, general and optimal solution. In Proceedings of the IEEE International Conference on Computer Vision, pp. 2344–2351, 2013.